OpenStudy (anonymous):

Solve for k: 1/10k - 1/15= -1/6 what is k

OpenStudy (mathlegend):

\[\frac{ 1 }{ 10 }k - \frac{ 1 }{ 15 } = -\frac{ 1 }{ 6 }\] So lets move the 1/15 to the other side.

OpenStudy (mathlegend):

@monica352?